OpenStudy (anonymous):

Factor completely: −5k2 + 20k − 30 −1(5k2 − 20k + 30) −5(k2 − 4k + 6) −5k(k2 − 4k + 6) −5(k2 + 4k − 6)

OpenStudy (callisto):

−5k^2 + 20k − 30 = -5k^2 + 4(5)k - 5(6) What is the common factor for these three terms?

OpenStudy (anonymous):

5 isnt it

OpenStudy (anonymous):

5 is it

OpenStudy (callisto):

Yes. You can also take out one more... Can you try?

OpenStudy (anonymous):

i think 5 is the only number that can go into all three terms other than 1 that is

OpenStudy (callisto):

You can also take out -1 as the common factor... Now, take out the common factors and group the rest of the terms, can you?

OpenStudy (anonymous):

−1(5k2 − 20k + 30) i think thats it

OpenStudy (callisto):

Not really. Didn't you mention 5 is also a common factor?

OpenStudy (anonymous):

im not sure

OpenStudy (callisto):

−5k^2 + 20k − 30 = -5k^2 + 4(5)k - 5(6) = -5k^2 - 4(-5)k - 5(6) Now, can you observe what is the common factor?

OpenStudy (anonymous):

give me a sec

OpenStudy (anonymous):

not really still dont quite understand

OpenStudy (callisto):

what are the factors for the first term?